OpenStudy (anonymous):

A set of equations is given below: Equation A: u = 4v + 1 Equation B: u = 3v + 5 Which of the following is a step that can be used to find the solution to the set of equations?

OpenStudy (anonymous):

A. 4v + 1 = 3v B. 4v = 3v + 5 C. 4v + 5 = 3v + 1 D. 4v + 1 = 3v + 5

OpenStudy (usukidoll):

hint: substitute one of the u = into one of the equations.

OpenStudy (anonymous):

b?

OpenStudy (usukidoll):

no... I'll give an example Equation A: x = 4y+1 Equation B x = 2y+2 so one of the steps is to substitute one of the equations into the other . So I'll take my equation A which is x = 4y+1 and place it to equation B so I should have 4y+1=2y+2

OpenStudy (anonymous):

Oh. Then it's D?

OpenStudy (usukidoll):

yes :)

OpenStudy (anonymous):

Thank you

OpenStudy (usukidoll):

yes :)